OpenStudy (anonymous):

∫(2x+y^2(x))dx+(2xy+3y)dy where C is bounded by y=x^2+1, x+y=13 and x=1 with positive orientation

OpenStudy (anonymous):

@rafeakeejr @LivvieLouTattoo @nobodyepic @nikosis01 @Amykins @iAd4m @Clueless_math @idellasosuper @idellasosuper @unlockyourfuture @Cody17 @wizfan @Can_I_Lick_Your_Face? @rafeakeejr @marissagatorgirl @sumanta @SODAPOP @sgayton27 @jacobwat @MissLovelyPearlOyster @animal_lover36 @xuvaze @Stephanizer @lkay @momoluvgd @momoluvgd @sheelo_mughal

OpenStudy (anonymous):

What ur question?

OpenStudy (anonymous):

post it as equation so that it will be easy for others to help u there is an option below the reply space

OpenStudy (anonymous):

\[\int\limits_{C}^{}(2x+y^2x)dx+(2xy+3y)dy\]

OpenStudy (anonymous):

is it algebra or other if its other then im sorry i dont know

OpenStudy (anonymous):

it is calculus 2 :(

OpenStudy (anonymous):

i've not studied it yet

OpenStudy (anonymous):

thanks anyway

OpenStudy (phi):

For each section of the curve, you must express x and y (and dx and dy) in terms of a single variable. For example, along the curve defined by y= x^2 + 1 (and dy = 2x dx) replace y with x^2+1 and dy with 2x dx in your expression you will get an expression in terms of x and dx integrate, using limits of x=1 to 3 (see attached)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\int\limits_{?}^{?}(2x-2xy)dxdy\] i think it is something like this but i need bounds

OpenStudy (phi):

A line integral along closed curve C can be evaluated using Green's Thm. if you make the inner integral dy, you would go from the lower curve y=x^2+1 to the upper curve y= 13-x the outer integral over dx is 1 to 3

OpenStudy (anonymous):

\[\int\limits_{1}^{3}\int\limits_{2}^{12}(-2yx+2y)dydx\]

OpenStudy (phi):

no, the y limits are a function of x y=x^2+1 to y= 13-x

OpenStudy (anonymous):

\[\int\limits_{1}^{3}\int\limits_{x^2+1}^{13-x}(-2yx+2y)dydx\]

OpenStudy (phi):

yes

OpenStudy (anonymous):

thanks